Monday 03.20 Jenn/Edgar
Manager Communication: Pretty quiet day for both the bars today. All the theaters were light in sales and neither bar saw a lot of foot traffic today. EA
BAR Staff Observations: Ashley and Eliel were able to get both bars restocked and reset after the weekend, with all frozies in great shape. Ashley was able to deepclean the shelves and the glassware mats on the backbar at Trees during the first round. JH
Openers were out sooner rather than later since sales levels were low and closers were able to handle the bars with no incident. No traffic after the last round of Scream let out so it was an easy close for Miguel. EA
Staff Meeting Agenda notes: A reminder for bartenders to make sure they're doing a walk through in the Mezz before doors. JH
Ordering/Receiving/Shipping: Big Geyser & Stumptown arriving tomorrow. JH
Staff Meeting Agenda notes: A reminder for bartenders to make sure they're doing a walk through in the Mezz before doors. JH
Ordering/Receiving/Shipping: Big Geyser & Stumptown arriving tomorrow. JH
Ordered Smoked Chili Bitters. Should be in on Friday at earliest, but we have enough to start the John Wick special. EA
Maintenance: The area where the coil meets the fitting on the back of the lobby coffee machine is now leaking. I have added this to the log and emailed Juan regarding the matter to see if this is something that can get fixed in house before I default back to the Ciro, our Fetco guy. EA
Maintenance: The area where the coil meets the fitting on the back of the lobby coffee machine is now leaking. I have added this to the log and emailed Juan regarding the matter to see if this is something that can get fixed in house before I default back to the Ciro, our Fetco guy. EA
Comments
Post a Comment